# Embedding-Based Context-Aware Reranker

### v1

@@ 5 Conclusion and Discussion
For reproducibility, we provide implementation details of our method in Section 4.3 and those of the baseline methods in Appendix A.2. The full codebase will be released publicly upon acceptance of this paper.
@@ Appendix A Appendix
A.1 Supplemental MRR@10 Results
@@
A.2 Supplemental Implementation Details
For monoBERT and monoT5 trained on ConTEB, we implement these two models from scratch. For training, we use the Adam optimizer (Kingma & Ba, 2015) with a fixed learning rate of $2\times 10^{-5}$ and $5\times 10^{-5}$, respectively, as suggested in their original implementation details (Nogueira et al., 2019; 2020). These models are trained for up to $20$ epochs, with early stopping based on validation loss using a patience of $5$ epochs. We set the batch size to $32$. All other baseline implementations are based on public code releases: monoT5 (MS-MARCO), duoT5, RankVicuna, RankZephyr, FirstMistral, and LiT5Distill are from the RankLLM library (Sharifymoghaddam et al., 2025) [^1] https://github.com/castorini/rank_llm. Following ICR (Chen et al., 2025), we use open-source models Mistral (Jiang et al., 2023) [^2] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2 and Llama-3.1 (et al., 2024) [^3] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct for the RankGPT method, instead of GPT3.53.5/44 due to cost constraints. FIRST [^4] https://github.com/gangiswag/llm-reranker, PE-Rank [^5] https://github.com/liuqi6777/pe_rank, RankGPT and ICR [^6] https://github.com/OSU-NLP-Group/In-Context-Reranking follow their respective original implementations. All experiments are run on a single A100100 GPU.
A.3 Performance on MS-MARCO V2
@@
A.4 Impact of Retrieved Candidate Size

### v2

@@ 5 Conclusion and Discussion
For reproducibility, we provide implementation details of our method in Section 4.3 and those of the baseline methods in Appendix A.3. The full codebase is released publicly at https://github.com/BorealisAI/EBCAR.
LLM Usage
We employed large language models exclusively to assist with polishing the writing of this paper. Specifically, LLMs were used to refine wording and verify grammar for clarity and readability. No aspects of method design, implementation, or experimental analysis involved LLM assistance.
@@ Appendix A Appendix
A.1 Training with Updated Query Embedding
Figure 4: Training loss comparison between using an updated query embedding (left) and a fixed query embedding (right) in Eq. 2. Updating the query embedding leads to large and unstable loss spikes, while keeping it fixed produces smooth and monotonic convergence. This supports our design choice to use a fixed query embedding.
To examine the effect of fixing the query embedding in the contrastive objective in Eq. 2, we compare two variants of EBCAR: one that uses the fixed input query embedding $q$, and another that uses the updated query representation produced by the encoder to compute the loss. As shown in Figure 4, training becomes highly unstable when the query embedding is updated. Loss values oscillate drastically and occasionally explode after a few epochs. In contrast, the fixed-query variant converges smoothly and yields stable optimization. We attribute this instability to context drift, where the query representation becomes entangled with passage-specific context, thus breaking its role as a consistent semantic anchor across candidate sets. These results empirically validate our design choice of using the unmodified query embedding in Eq. 2.
A.2 Supplemental MRR@10 Results
@@
A.3 Supplemental Implementation Details
For monoBERT and monoT5 trained on ConTEB, we implement these two models from scratch. For training, we use the Adam optimizer (Kingma and Ba, 2015) with a fixed learning rate of $2\times 10^{-5}$ and $5\times 10^{-5}$, respectively, as suggested in their original implementation details (Nogueira et al., 2019; 2020). These models are trained for up to $20$ epochs, with early stopping based on validation loss using a patience of $5$ epochs. We set the batch size to $32$. All other baseline implementations are based on public code releases: monoT5 (MS-MARCO), duoT5, RankVicuna, RankZephyr, FirstMistral, and LiT5Distill are from the RankLLM library (Sharifymoghaddam et al., 2025) [^2] https://github.com/castorini/rank_llm. Following ICR (Chen et al., 2025), we use open-source models Mistral (Jiang et al., 2023) [^3] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2 and Llama-3.1 (et al., 2024) [^4] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct for the RankGPT method, instead of GPT3.53.5/44 due to cost constraints. FIRST [^5] https://github.com/gangiswag/llm-reranker, PE-Rank [^6] https://github.com/liuqi6777/pe_rank, RankGPT and ICR [^7] https://github.com/OSU-NLP-Group/In-Context-Reranking follow their respective original implementations. All experiments are run on a single A100100 GPU.
A.4 Examples of Football and Insurance
Example 1 (Football). Query: What was Andrew Henry Robertson’s first goal for Hull City? Gold passage: Despite the departure of several other first-team players, he chose to remain at City. His inaugural goal for the club was scored on 3 November 2015 in an away match against Brentford, where he initiated the scoring in a 2–0 victory that placed Hull at the top of the Championship table on goal difference. He participated in the 2016 Championship play-off final against Sheffield Wednesday, which Hull won 1–0 to achieve promotion to the Premier League. However, the team lasted only one season in the top division before being relegated again.
All pronouns in this dataset are anonymized, so identifying that ”he” refers to ”Andrew Henry Robertson” requires reasoning across multiple passages mentioning the player and the club. EBCAR’s hybrid attention enables such cross-passage coreference resolution effectively.
Example 2 (Insurance). Query: What is the amount of (Re)insurance GWP in million in Austria? Gold passage: {’General data’: ’— Amounts — Share total EEA —
— Population (in 1000) — 8,979 — 2.00% —
— (Re)insurance GWP (in million) — 20,815.873 — 1.5% —
— Number of (re)insurance undertakings — 33 — 1.9% —
— Number of registered insurance intermediaries — 17,999 — 2.1% —’
Without positional or structural cues, it is difficult to associate this numerical information with the correct country (Austria), which only appears in the section title. EBCAR’s positional encodings provide this structural grounding, allowing accurate passage ranking.
A.5 Hybrid Attention Analysis
To examine how the hybrid attention mechanism routes information, we conduct a qualitative study across all datasets. For each dataset, we randomly sample $20$ test queries and, for every example, render an aggregated attention heatmap averaged over heads. We focus our analysis on the final layer (Layer 15) where the model’s decision is consolidated. Figure 6 shows the shared full attention at Layer 15. Although weights vary across cells, attention is broadly distributed across passages, indicating global context exchange among all candidates. In contrast, Figure 7 shows the dedicated masked attention at Layer 15: information flow is sharply concentrated within passages that originate from the same document (block-diagonal patterns) and between the query and those passages. While the figures present one representative case, we observe the same qualitative pattern consistently across the 20 sampled examples in every dataset: the shared full attention supports global cross-passage reconciliation, whereas the dedicated masked attention strengthens intra-document consolidation. This matches EBCAR’s design intent and explains the complementary gains seen in the ablations.
A.6 Performance on MS-MARCO V2
@@
A.7 Impact of Retrieved Candidate Size
@@
Figure 6: Shared full attention at Layer 15, averaged over all heads. The green box highlights the gold passage. Attention is more broadly distributed across passages, indicating global context exchange among all candidates.
Figure 7: Dedicated masked attention at Layer 15, averaged over all heads. The green box highlights the gold passage. Attention concentrates within passages from the same document (block-diagonal structure) and between the query and those passages, reflecting document-aware locality.